OpenStudy (anonymous):

Again. -.- 4x-y=2 5x-2y=-2

OpenStudy (anonymous):

y = 4x -2 5x -2(4x - 2) = -2 5x +4 - 8x = -2 -3x = -6 x = 2

OpenStudy (anonymous):

y = 4x -2 y = 4(2) -2 y = 8 -2 y = 6

OpenStudy (anonymous):

4x-y=2 [1] 5x-2y=-2 [2] make y the subject in equation [1] y = 2-4x substitute into [2] 5x-2(2-4x )=-2 expand 5x-4+8x =-2 13x -4 = -2 13x = 2 x = 2/13 Subsstitute for x in [1] 8/13-y=2 y = -18/13

OpenStudy (anonymous):

You forgot that there is a negative sign in y.

OpenStudy (anonymous):

What? His answer is wrong. Why the medal? :( Look at my solution.